OpenStudy (rayford):

dentify the discontinuities, graph, and find the zeros of each rational function. f(x)=2x^2+5x-12/x+4

OpenStudy (anonymous):

we know the discontinuity is where \(x+4=0\) i.e. \(x=4\)

OpenStudy (anonymous):

zeros will be where \[2x^2+5x-12=0\] maybe it factors

OpenStudy (anonymous):

i suck at factoring, but this one is \((2x-3)(x+4)\)zeros are easy enough to find what else do you need?

OpenStudy (rayford):

oh and a graph i believe.

OpenStudy (rayford):

did he answer it???

OpenStudy (mathmate):

Yes, he gave you the complete answer, like last time. The discontinuity is at x+4=0 or x=-4. I.e. the function is undefined at x=-4. The remaining points form a straight line where y=2x-3. The problem is similar to the previous one.

OpenStudy (rayford):

well is is -4 or 4 ?

OpenStudy (rayford):

it's 4...

OpenStudy (mathmate):

It's x+4=0, which means that the discontinuity is at x=-4.